An online stock trading firm has a fixed cost of $595 and a variable cost of $12x, where x is the number of clients that subscribe to the firm's trading service. If the firm has 120 clients, what is the lowest price it can charge each client without pushing total revenue below cost?

Group of answer choices

$14.9

$15.9

$17

$20
Answer #1

Total cost for 120 clients = 595 + 12 * 120 = 2035

Minimum price to be charged = 2035 / 120 = 16.958 ~ 17

Third option is correct